OpenStudy (anonymous):

A container has 50 electronic components, of which 10 are defective. If 8 components are drawn at random from the container, what is the probability that exactly 3 of them are defective? A-0.147 B-0.203 C-0.300 D-0.375 E-0.750

OpenStudy (kropot72):

The probability of a component being defective is 10/50 = 0.2. \[P( out\ of\ 8\ defective)=8C3\times(0.2)^{3}\times(0.8)^{5}=you\ can\ calculate\]